OpenStudy (daisy.xoxo619):

G: <1 COMP<2 <3 COMP<2 P: QT PARALLEL RS WRITE 2 COLUMN PROOF

OpenStudy (daisy.xoxo619):



OpenStudy (daisy.xoxo619):

@jim_thompson5910 stuck again please help

jimthompson5910 (jim_thompson5910):

show me what you have so far

OpenStudy (daisy.xoxo619):

1. <1 COMP <2 1. GIVEN 2. <3 COMP <2 2. GIVEN 3. <1 & <2 RT ANGLE 3. DEG OF RT ANGLES 4. <3 RT ANGLE 4. DEF OF RIGHT ANGLES

jimthompson5910 (jim_thompson5910):

how do you know that " <1 & <2 RT ANGLE" ?

OpenStudy (daisy.xoxo619):

ARENT 2 COMP ANGLES EACH 90 DEGREES SO IT FORMS RIGHT ANGLES

jimthompson5910 (jim_thompson5910):

they aren't 90 degrees on their own they are complementary angles, so they add to 90

OpenStudy (daisy.xoxo619):

oh ok

jimthompson5910 (jim_thompson5910):

so say m < 1 = x m < 2 = y m < 3 = z

jimthompson5910 (jim_thompson5910):

angles 1 and 2 are complementary, so, x+y = 90

jimthompson5910 (jim_thompson5910):

2 and 3 are complementary y+z = 90

OpenStudy (daisy.xoxo619):

so what would my reasons be for : m < 1 = x m < 2 = y m < 3 = z and x+y = 90 and y+z = 90

jimthompson5910 (jim_thompson5910):

the first three lines are simply definitions or assignments

jimthompson5910 (jim_thompson5910):

we're letting angle 1 be x, 2 be y, 3 be z the last two lines are basically true because of the definition of complementary.

OpenStudy (daisy.xoxo619):

so now i think the last step would be the prove but can't figure out the reason

jimthompson5910 (jim_thompson5910):

hint: x+y = 90 z+y = 90 what can you say about x and z?

OpenStudy (daisy.xoxo619):

there corresponding angles?

jimthompson5910 (jim_thompson5910):

what else

OpenStudy (daisy.xoxo619):

there cong.

jimthompson5910 (jim_thompson5910):

correct, so x = z meaning that angle 1 is congruent to angle 3

jimthompson5910 (jim_thompson5910):

What's the next step?

OpenStudy (daisy.xoxo619):

im not sure

jimthompson5910 (jim_thompson5910):

1 and 3 are corresponding angles

jimthompson5910 (jim_thompson5910):

they are congruent so that must mean these two lines are parallel 

OpenStudy (daisy.xoxo619):

ok so can i just say if 2 angles are cong then the 2 lines are parallel

jimthompson5910 (jim_thompson5910):

yes because of the corresponding angles converse

OpenStudy (daisy.xoxo619):

ok thank you:)

jimthompson5910 (jim_thompson5910):

you're welcome